OpenStudy (maeganr):

Point-Slope Form: y+3=2(x+1)

OpenStudy (anonymous):

What do you need help with?

OpenStudy (maeganr):

ok. um. how do you solve and graph it? ill show you what ive already done.

OpenStudy (maeganr):

y+3=-2(x+1) y+3=-2x+2 -3 -3 y=-2x... Thats where im stuck.

OpenStudy (maeganr):

...

OpenStudy (anonymous):

To find the points so you can graph it, just plug 0 in for x or y: y+3=2(x+1) 0+3=2(x+1) 3=2x+1 2=2x 1=x Here's a point: (1, 0) y+3=2(x+1) y+3=2(0+1) y+3=2 y= -1 Heres another point: (0, -1) You can plug any number in for x or y to find the a point. :o)

OpenStudy (anonymous):

Also, is it 2(x+1) or -2(x+1)?

OpenStudy (maeganr):

y+3=-2(x+1)

OpenStudy (anonymous):

Okay, then ignore that last comment lol. y+3=-2(x+1) 0+3=-2x-2 3=-2x-2 5=-2x -2.5=x (-2.5, 0) y+3=-2(0+1) y+3=-2 y=-5 (0, -5)

OpenStudy (maeganr):

oh. |:

OpenStudy (anonymous):

You can do this and find a few different points and then you can graph them. :o)

OpenStudy (maeganr):

ok

OpenStudy (maeganr):

can you help me with y+1=(x+5) also?

OpenStudy (anonymous):

@MaeganR didn't word the question too well. She said "solve" as in put it in point intercept form: y=mx+b

OpenStudy (maeganr):

y+1=-3/5(x+5) oops.

OpenStudy (anonymous):

MaeganR. Expand/Distribute the LHS (Left Hand Side)

OpenStudy (maeganr):

i should really write that down..